OpenStudy (anonymous):

find the solutions to the system. y = x^2 - 3x - 1 y=8x - 1 a. (0,-1) and (11,388) b. (0,01) and (11,87) c. (-1,0) and (87,11) d. no solutions

OpenStudy (anonymous):

do you have an idea of what to do?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so let's use substitution... since the y's have to be the same value (as do the x's) we can equate them and get an equation with only x's... 8x-1 = y = x^2 - 3x - 1 => 8x-1 = x^2 - 3x - 1 now you can move everything to one side and solve the quadratic

OpenStudy (anonymous):

do you know how to solve?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

subtract (8x-1) from both sides... what do you get?

OpenStudy (anonymous):

what is x